Grade 12th passInorganic Chemistry

Number of molecules of CH4 in 80.0 g of it.....??????????

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer:

ch4=80.0g

molecules =12+4=16

=80/16

=5

ch4 molecules =5

Thanks